# Diffusion Meets DAgger: Supercharging Eye-in-hand Imitation Learning

### v1

@@ V Discussion
DMD was designed to improve the sample efficiency of eye-in-hand imitation learning. Compared to traditional behavior cloning, DMD shows significant improvement in experiments involving non-prehensile pushing across 2 objects. Remarkably, DMD provides effective augmenting strategies using generated images and enabled learned policies to achieve 80% success rate from as few as 8 demonstrations. DMD not only performs better than conventional augmentation schemes used in robotics but also shows improved metrics when combined with them.
Our experiments revealed the benefits of the different parts of DMD over past approaches. The use of diffusion models to synthesize images allows DMD to generate augmenting images for manipulation tasks that involve non-static scene. Prior work [85] used NeRFs instead of diffusion models, restricting generation to samples from pre-grasp part of the trajectories. We showcase the advantages of diffusion models over NeRFs through qualitative results in Figure 8 and real robot experiments in Section IV-D3. We also tested which type of data: task data, play data, or a combination of both, should be used for training the diffusion model. We found all versions to outperform behavior cloning. This suggests two practically useful results. First, DMD can directly be used to improve a task policy without the need to collect any additional data, i.e. the task data itself can be used to finetune a novel-view synthesis diffusion model to augment the task dataset to improve performance. Second we can benefit from play data, which maybe simpler to collect than task data. To foster replication and future work, we will make code, data, and models publicly available.
@@ Acknowledgments
We give special thanks to Kevin Zhang for 3D printing the attachment for us. This material is based upon work supported by the USDA/NSF AIFARMS National AI Institute (USDA #2020-67021-32799) and NSF (IIS-2007035).

### v2

@@ V Discussion
We developed DMD, a data creation framework that improves the sample efficiency of eye-in-hand imitation learning. DMD uses a diffusion model to synthesize out-of-distribution views and assigns them corrective labels. Additional use of these out-of-distribution views for policy training leads to more performant policies. DMD leads to large benefits over traditional behavior cloning across 4 diverse tasks: pushing, stacking, pouring, and hanging a shirt. DMD attains 80% success rate from as few as 8 demonstrations in pushing and reaches an average of 92% success rate across 5 different cups in stacking. When pouring, it transfers coffee beans successfully 80% of the time; finally, it can slide a shirt off the table and put it on the rack with 90% success rate. Together these tasks test DMD on different aspects: 3DoF action space (pushing, stacking), 6DoF action spaces (pouring, hanging a shirt), generalization to new objects (stacking), and precision in reaching goal locations (pouring, hanging a shirt).
Our experiments reveal the benefits of the various parts of DMD over past approaches. The use of diffusion models to synthesize images allows DMD to generate augmenting images for manipulation tasks involving a non-static scene. Prior work [86] used NeRFs instead of diffusion models, restricting generation to samples from only the pre-grasp part of the demonstration. We showcase the advantages of diffusion models over NeRFs through qualitative results in Figure 9 and real robot experiments in Section IV-A4.
We also tested which type of data: task data, play data, or a combination of both, should be used for training the diffusion model. We found all versions to outperform behavior cloning (Section IV-A4, Table III, and Table IV). This suggests two practically useful results. First, DMD can be directly used to improve a task policy, i.e. the task data itself can be used to finetune a novel-view synthesis diffusion model, augmenting the task dataset to improve performance. Second, we can utilize play data, which may be simpler to collect than task data.
DMD also has several limitations. DMD assumes that the generated states are recoverable, i.e. there exists an action that can return to the expert distribution. This is not always true, for example, when the robot rotates the cup at the wrong location and the coffee beans spill. However, DMD can still be useful if it can bring the system back to in-distribution states before the system spirals into irrecoverable state.
Possible future directions include: learning diffusion models that can predict highly discontinuous object dynamics, generating out-of-distribution states of non-visual modalities (e.g. forces), and finding ways to assign action labels in situations where they cannot be derived from relative camera transformation. Code, data and models are publicly available on the project website.
@@ Acknowledgments
We give special thanks to Kevin Zhang for 3D printing the attachment for us. This material is based upon work supported by the USDA-NIFA AIFARMS National AI Institute (USDA #2020-67021-32799) and NSF (IIS-2007035).